# Chibueze Peace Obioma    Youcheng Sun    Mustafa A. Mustafa

### v2

@@ 1 Introduction
Federated learning (FL) enables collaborative model training across distributed clients without centralizing sensitive data [mcmahan2017communication]. This architecture has found adoption in privacy-critical domains including healthcare diagnostics [nguyen2022federated, silva2019federated], financial fraud detection [li2022survey], and mobile keyboard prediction [kairouz2021advances]. However, the server’s inability to inspect local data or verify computation integrity creates opportunities for Byzantine adversaries to inject malicious updates that compromise the global model.
@@
Most existing defenses operate from a premise that malicious behavior manifests as statistical anomalies [cao2021fltrust, blanchard2017machine]. Byzantine clients are expected to exhibit deviations from the benign majority through gradient magnitude outliers, divergent parameter directions, or aberrant update patterns, leading detection methods to flag clients whose statistical signatures fall outside expected bounds. While some attacks [bagdasaryan2020backdoor, xie2020dba] do manifest as statistical outliers, this anomaly-centric perspective captures one dimension of the threat surface but overlooks a fundamental constraint: an attacker embedding a backdoor trigger must ensure reliable misclassification across hundreds of training rounds, despite model parameter drift, aggregation with benign updates, and stochastic client sampling. This reliability demands behavioral consistency. This consistency requirement creates detection surfaces beyond traditional anomaly metrics.
The backdoor threat landscape exhibits significant diversity in how attacks manifest this consistency constraint. Some attacks [gu2017badnets, wang2020attack] prioritize reliability over stealth, accepting higher consistency signatures to ensure trigger activation. Others [xie2020dba, yang2024distributed] distribute the backdoor across colluding clients, fragmenting the consistency pattern but creating coordinated behavioral signatures. Adaptive attacks [fang2020local, zhang2023a3fl, wen2022thinking] calibrate poisoning intensity across rounds, trading reliability for reduced detectability, while scaling attacks [bagdasaryan2020backdoor] inject high-magnitude updates to dominate aggregation. This diversity reveals that different attack strategies ensure consistency through distinct mechanisms, each creating characteristic signatures across orthogonal dimensions. We empirically observe this pattern across multiple attack types (see Section 4), and this observation motivates taking a new detection approach targeting consistency patterns rather than anomalies alone.
@@ 2.2 Backdoor Attacks in Federated Learning
Backdoor attacks in FL exhibit diverse sophistication levels. Model replacement attacks [bagdasaryan2020backdoor] scale malicious updates by factor $\lambda=K/|M|$ where $K$ is total clients and $|M|$ is malicious count, aiming to dominate aggregation through magnitude manipulation. Distributed backdoor attacks (DBA) [xie2020dba] decompose triggers across colluding clients, making individual updates appear benign while collectively embedding backdoors. Edge-case attacks [wang2020attack] target rare inputs on distribution tails, exploiting natural class confusion patterns for highly stealthy backdoors. Neurotoxin [zhang2022neurotoxin] achieves durability through momentum-guided parameter selection, slowly-changing parameters identified through gradient momentum analysis, ensuring backdoor persistence across hundreds of rounds. IBA [nguyen2023iba] combines learnable trigger optimization with adversarial perturbations as triggers, making detection via visual inspection ineffective. Recent work explores stealthy attacks against personalized FL [lyu2024lurking], data-free backdoor injection [li2024darkfed], and distributed attacks on graph learning [yang2024distributed]. These adaptive attacks specifically design updates to satisfy defense constraints while preserving malicious functionality [wang2020attack].
@@ 2.3 Defenses Against Backdoor Attacks
Pre-aggregation defenses. These methods filter malicious updates before aggregation by analyzing individual client contributions. FoolsGold [fung2020foolsgold] exploit learning history to identify clients with similar gradient patterns, reducing weights for coordinated attackers. FLAME [nguyen2022flame] employs HDBSCAN clustering on client representations, excluding minority clusters as potentially malicious. DeepSight [rieger2022deepsight] applies spectral clustering on representation covariance matrices, and carries out deep model inspection to analyze internal representations for anomalous patterns. Trust bootstrapping methods like FLTrust [cao2021fltrust] maintain a small root dataset on the server, computing cosine similarity between client updates and trusted root updates to filter anomalous contributions. Other recent works that falls under this category include FedAvgCKA [walter2024exploiting], frequency analysis method [fereidooni2024freqfed], out-of-distribution data detection [li2024backdoorindicator], federated detection [rieger2024crowdguard], and certified robustness [xie2021crfl].
@@ 2.4 Critical Vulnerabilities in Existing Defenses
First, existing defenses rely on static detection assumptions, expecting malicious updates to exhibit consistent deviations from benign patterns under fixed metrics. Byzantine-robust aggregation rules such as MKrum [blanchard2017machine] and Bulyan [mhamdi2018hidden] operate through distance-based outlier detection and geometric median computation, achieving provable robustness against untargeted poisoning. However, sophisticated attackers normalize update magnitudes while preserving backdoor functionality through careful parameter crafting, violating the fixed-deviation assumption. Defenses like FLTrust [cao2021fltrust], maintain a root dataset for trust bootstrapping, computing cosine similarity between client updates and trusted updates to filter anomalous contributions. Yet adaptive attackers exploit these single-metric approaches through coordinate-wise optimization, maximizing similarity scores while maintaining backdoor effectiveness [bagdasaryan2020backdoor]. These defenses implicitly assume that malicious representations maintain fixed geometric relationships to benign ones, an assumption that collapses under adaptive manipulation.
@@ 3.1 Attacker Model
Attacker limitations. We assume the attacker cannot compromise the server, cannot control the majority of clients, and cannot observe other clients’ local data or models. However, we explicitly allow malicious clients to share and coordinate their updates with each other, as observed in distributed backdoor attacks [xie2020dba]. These assumptions are standard in FL security research [kairouz2021advances].
@@ 4.1 Observations and Design Choices
We investigate how backdoor injection manifests across multiple behavioral dimensions through controlled synthetic experiments. Following established backdoor methodologies [bagdasaryan2020backdoor, wang2020attack, xie2020dba], we constructed FL scenarios where Byzantine clients embed various trigger patterns into local training data while maintaining high clean accuracy to evade performance-based detection.
@@
Beyond variance suppression, we observe elevated inter-client behavioral coordination in scenarios where multiple malicious clients participate (e.g., distributed backdoor attacks [xie2020dba]). Fig. 4 demonstrates that colluding attackers, despite training on different local data partitions, exhibit systematically higher pairwise parameter similarity compared to benign client pairs. This coordination emerges because colluding attackers optimize identical backdoor objectives, mapping trigger pattern $\tau$ to target label $y_{t}$ regardless of local data differences [fang2020local]. While benign clients’ parameter trajectories diverge due to heterogeneous task distributions, malicious clients converge toward shared trigger-aligned representations. Critically, this pattern persists even when individual malicious updates are normalized to appear benign in magnitude or direction. The coordination signature manifests in the upper tail of mutual similarity metrics.
@@ 4.3.4 Adaptive Aggregation
Defense No Attack BadNet [gu2017badnets] Blended [chen2017targeted] Edge-case [wang2020attack] IBA [nguyen2023iba] Neurotoxin [zhang2022neurotoxin] Chameleon [dai2023chameleon] Avg. MA Avg. BA
@@ 5.1 Experimental Setup
Datasets and models. We evaluate FeRA on four widely-used image classification benchmarks: CIFAR-10, CIFAR-100 [krizhevsky2009learning], Tiny-ImageNet, and German Traffic Sign Recognition Benchmark (GTSRB) [GTSRB2012Zenodo] dataset. For model architectures, we employ ResNet-18 [He2016DeepResidual] as our primary architecture, with additional evaluation on ResNet-34 [He2016DeepResidual] and VGG-16 [Simonyan2015VeryDeep] to demonstrate architecture-agnostic effectiveness.
@@
We consider both IID and non-IID data partitions. For the IID case, training data are evenly distributed across clients. For the non-IID case, we follow the standard Dirichlet procedure [hsu2019measuring, huang2023multimetrics, huang2024lockdown], drawing each client’s data proportion from $\text{Dir}(\alpha)$ with default $\alpha=0.5$. Local training uses SGD with server learning rate $\eta=0.5$, client learning rate $\alpha=0.1$, and two local epochs per round. As is common in FL, we use dataset-specific pretraining before enabling attacks: CIFAR-10 is pretrained for 2000 rounds, CIFAR-100 for 1000, and Tiny-ImageNet for 500; GTSRB and MNIST variants are trained from scratch, with attacks starting after 50 and 5 rounds, respectively. Unless otherwise specified, attacks are then run for 100 rounds. The server aggregates updates using FedAvg [mcmahan2017communication].
@@
Attack scenarios. We evaluate against eight attack variants: BadNet [gu2017badnets], Blended [chen2017targeted], Edge-case [wang2020attack], DBA [xie2020dba], Neurotoxin [zhang2022neurotoxin],Chameleon [dai2023chameleon], IBA [nguyen2023iba], A3FL [zhang2023a3fl], and an adaptive attack. In our ablation studies we also evaluate against Anticipate [wen2022thinking] attack which by principle of its design can evade some of our defense filter. During training, attackers may control single or multiple clients, initiating poisoning from the poisoning starting round as described initially.
Baseline defenses. We compare FeRA with the non-robust baseline FedAvg and against state-of-the-art defenses: Multi-Krum [blanchard2017machine], FoolsGold [fung2020foolsgold], FLAME [nguyen2022flame], FLTrust [cao2021fltrust], RLR [ozdayi2021defending], Lockdown [huang2024lockdown] and Deepsight [rieger2022deepsight].
@@ 5.4 Resilience Under Data Heterogeneity
Pattern [gu2017badnets] 72.04 1.85 85.47 1.16 91.26 1.53 91.35 1.14 91.84 1.36 86.39 1.41
Pixel [gu2017badnets] 82.47 5.63 85.05 0.74 90.05 0.94 90.26 1.63 91.85 1.26 87.94 2.04
BadNet [gu2017badnets] 80.16 15.74 84.36 0.84 91.63 1.26 90.63 1.74 91.74 1.36 87.70 4.19
@@ 5.5 Cross-Silo Federated Learning
Lockdown [huang2024lockdown] 88.14 / 6.35 81.36 / 10.84 84.75 8.60

### v3

@@ 1 Introduction
Federated learning (FL) enables collaborative model training across distributed clients without centralizing sensitive data [mcmahan2017communication]. This architecture has found adoption in privacy-critical domains including healthcare diagnostics [nguyen2022federated, silva2019federated], financial fraud detection [li2021survey], and mobile keyboard prediction [kairouz2021advances]. However, the server’s inability to inspect local data or verify computation integrity creates opportunities for Byzantine adversaries to inject malicious updates that compromise the global model.
@@
Most existing defenses operate from a premise that malicious behavior manifests as statistical anomalies [cao2020fltrust, blanchard2017machine]. Byzantine clients are expected to exhibit deviations from the benign majority through gradient magnitude outliers, divergent parameter directions, or aberrant update patterns, leading detection methods to flag clients whose statistical signatures fall outside expected bounds. While some attacks [bagdasaryan2020backdoor, xie2019dba] do manifest as statistical outliers, this anomaly-centric perspective captures one dimension of the threat surface but overlooks a fundamental constraint: an attacker embedding a backdoor trigger must ensure reliable misclassification across hundreds of training rounds, despite model parameter drift, aggregation with benign updates, and stochastic client sampling. This reliability demands behavioral consistency. This consistency requirement creates detection surfaces beyond traditional anomaly metrics.
The backdoor threat landscape exhibits significant diversity in how attacks manifest this consistency constraint. Some attacks [gu2019badnets, wang2020attack] prioritize reliability over stealth, accepting higher consistency signatures to ensure trigger activation. Others [xie2019dba, yang2024distributed] distribute the backdoor across colluding clients, fragmenting the consistency pattern but creating coordinated behavioral signatures. Adaptive attacks [fang2020local, zhang2023a3fl, wen2022thinking] calibrate poisoning intensity across rounds, trading reliability for reduced detectability, while scaling attacks [bagdasaryan2020backdoor] inject high-magnitude updates to dominate aggregation. This diversity reveals that different attack strategies ensure consistency through distinct mechanisms, each creating characteristic signatures across orthogonal dimensions. We empirically observe this pattern across multiple attack types (see Section 4), and this observation motivates taking a new detection approach targeting consistency patterns rather than anomalies alone.
@@ 2.2 Backdoor Attacks in Federated Learning
Backdoor attacks in FL exhibit diverse sophistication levels. Model replacement attacks [bagdasaryan2020backdoor] scale malicious updates by factor $\lambda=K/|M|$ where $K$ is total clients and $|M|$ is malicious count, aiming to dominate aggregation through magnitude manipulation. Distributed backdoor attacks (DBA) [xie2019dba] decompose triggers across colluding clients, making individual updates appear benign while collectively embedding backdoors. Edge-case attacks [wang2020attack] target rare inputs on distribution tails, exploiting natural class confusion patterns for highly stealthy backdoors. Neurotoxin [zhang2022neurotoxin] achieves durability through momentum-guided parameter selection, slowly-changing parameters identified through gradient momentum analysis, ensuring backdoor persistence across hundreds of rounds. IBA [nguyen2023iba] combines learnable trigger optimization with adversarial perturbations as triggers, making detection via visual inspection ineffective. Recent work explores stealthy attacks against personalized FL [lyu2024lurking], data-free backdoor injection [li2024darkfed], and distributed attacks on graph learning [yang2024distributed]. These adaptive attacks specifically design updates to satisfy defense constraints while preserving malicious functionality [wang2020attack].
@@ 2.3 Defenses Against Backdoor Attacks
Pre-aggregation defenses. These methods filter malicious updates before aggregation by analyzing individual client contributions. FoolsGold [fung2020foolsgold] exploit learning history to identify clients with similar gradient patterns, reducing weights for coordinated attackers. FLAME [nguyen2022flame] employs HDBSCAN clustering on client representations, excluding minority clusters as potentially malicious. DeepSight [rieger2022deepsight] applies spectral clustering on representation covariance matrices, and carries out deep model inspection to analyze internal representations for anomalous patterns. Trust bootstrapping methods like FLTrust [cao2020fltrust] maintain a small root dataset on the server, computing cosine similarity between client updates and trusted root updates to filter anomalous contributions. Other recent works that falls under this category include FedAvgCKA [walter2024exploiting], frequency analysis method [fereidooni2024freqfed], out-of-distribution data detection [li2024backdoorindicator], federated detection [rieger2024crowdguard], and certified robustness [xie2021crfl].
@@ 2.4 Critical Vulnerabilities in Existing Defenses
First, existing defenses rely on static detection assumptions, expecting malicious updates to exhibit consistent deviations from benign patterns under fixed metrics. Byzantine-robust aggregation rules such as MKrum [blanchard2017machine] and Bulyan [mhamdi2018hidden] operate through distance-based outlier detection and geometric median computation, achieving provable robustness against untargeted poisoning. However, sophisticated attackers normalize update magnitudes while preserving backdoor functionality through careful parameter crafting, violating the fixed-deviation assumption. Defenses like FLTrust [cao2020fltrust], maintain a root dataset for trust bootstrapping, computing cosine similarity between client updates and trusted updates to filter anomalous contributions. Yet adaptive attackers exploit these single-metric approaches through coordinate-wise optimization, maximizing similarity scores while maintaining backdoor effectiveness [bagdasaryan2020backdoor]. These defenses implicitly assume that malicious representations maintain fixed geometric relationships to benign ones, an assumption that collapses under adaptive manipulation.
@@ 3.1 Attacker Model
Attacker limitations. We assume the attacker cannot compromise the server, cannot control the majority of clients, and cannot observe other clients’ local data or models. However, we explicitly allow malicious clients to share and coordinate their updates with each other, as observed in distributed backdoor attacks [xie2019dba]. These assumptions are standard in FL security research [kairouz2021advances].
@@ 4.1 Observations and Design Choices
We investigate how backdoor injection manifests across multiple behavioral dimensions through controlled synthetic experiments. Following established backdoor methodologies [bagdasaryan2020backdoor, wang2020attack, xie2019dba], we constructed FL scenarios where Byzantine clients embed various trigger patterns into local training data while maintaining high clean accuracy to evade performance-based detection.
@@
Beyond variance suppression, we observe elevated inter-client behavioral coordination in scenarios where multiple malicious clients participate (e.g., distributed backdoor attacks [xie2019dba]). Fig. 4 demonstrates that colluding attackers, despite training on different local data partitions, exhibit systematically higher pairwise parameter similarity compared to benign client pairs. This coordination emerges because colluding attackers optimize identical backdoor objectives, mapping trigger pattern $\tau$ to target label $y_{t}$ regardless of local data differences [fang2020local]. While benign clients’ parameter trajectories diverge due to heterogeneous task distributions, malicious clients converge toward shared trigger-aligned representations. Critically, this pattern persists even when individual malicious updates are normalized to appear benign in magnitude or direction. The coordination signature manifests in the upper tail of mutual similarity metrics.
@@ 4.3.4 Adaptive Aggregation
Defense No Attack BadNet [gu2019badnets] Blended [chen2017targeted] Edge-case [wang2020attack] IBA [nguyen2023iba] Neurotoxin [zhang2022neurotoxin] Chameleon [dai2023chameleon] Avg. MA Avg. BA
@@ 5.1 Experimental Setup
Datasets and models. We evaluate FeRA on four widely-used image classification benchmarks: CIFAR-10, CIFAR-100 [krizhevsky2009learning], Tiny-ImageNet, and German Traffic Sign Recognition Benchmark (GTSRB) [GTSRB2012Zenodo] dataset. For model architectures, we employ ResNet-18 [He2016DeepResidual] as our primary architecture, with additional evaluation on ResNet-34 [He2016DeepResidual] and VGG-16 [Simonyan2014VeryDeep] to demonstrate architecture-agnostic effectiveness.
@@
We consider both IID and non-IID data partitions. For the IID case, training data are evenly distributed across clients. For the non-IID case, we follow the standard Dirichlet procedure [hsu2019measuring, huang2023multimetrics, huang2023lockdown], drawing each client’s data proportion from $\text{Dir}(\alpha)$ with default $\alpha=0.5$. Local training uses SGD with server learning rate $\eta=0.5$, client learning rate $\alpha=0.1$, and two local epochs per round. As is common in FL, we use dataset-specific pretraining before enabling attacks: CIFAR-10 is pretrained for 2000 rounds, CIFAR-100 for 1000, and Tiny-ImageNet for 500; GTSRB and MNIST variants are trained from scratch, with attacks starting after 50 and 5 rounds, respectively. Unless otherwise specified, attacks are then run for 100 rounds. The server aggregates updates using FedAvg [mcmahan2017communication].
@@
Attack scenarios. We evaluate against eight attack variants: BadNet [gu2019badnets], Blended [chen2017targeted], Edge-case [wang2020attack], DBA [xie2019dba], Neurotoxin [zhang2022neurotoxin],Chameleon [dai2023chameleon], IBA [nguyen2023iba], A3FL [zhang2023a3fl], and an adaptive attack. In our ablation studies we also evaluate against Anticipate [wen2022thinking] attack which by principle of its design can evade some of our defense filter. During training, attackers may control single or multiple clients, initiating poisoning from the poisoning starting round as described initially.
Baseline defenses. We compare FeRA with the non-robust baseline FedAvg and against state-of-the-art defenses: Multi-Krum [blanchard2017machine], FoolsGold [fung2020foolsgold], FLAME [nguyen2022flame], FLTrust [cao2020fltrust], RLR [ozdayi2021defending], Lockdown [huang2023lockdown] and Deepsight [rieger2022deepsight].
@@ 5.4 Resilience Under Data Heterogeneity
Pattern [gu2019badnets] 72.04 1.85 85.47 1.16 91.26 1.53 91.35 1.14 91.84 1.36 86.39 1.41
Pixel [gu2019badnets] 82.47 5.63 85.05 0.74 90.05 0.94 90.26 1.63 91.85 1.26 87.94 2.04
BadNet [gu2019badnets] 80.16 15.74 84.36 0.84 91.63 1.26 90.63 1.74 91.74 1.36 87.70 4.19
@@ 5.5 Cross-Silo Federated Learning
Lockdown [huang2023lockdown] 88.14 / 6.35 81.36 / 10.84 84.75 8.60